Using the data in the following table, estimate the:

a. Average return and volatility for each stock 

b. Correlation between these two stocks. 

c. The average return of stock A is %. 

Year201020112012201320142015
Stock A-13%20%9%-6%1%10%
Stock B21%33%45%-8%-7%33%
Answer #1
Stock A
Year AR ER = Avg. R (AR-ER) (AR-ER)^2
2010 -13 3.5 -16.5 272.25
2011 20 3.5 16.5 272.25
2012 9 3.5 5.5 30.25
2013 -6 3.5 -9.5 90.25
2014 1 3.5 -2.5 6.25
2015 10 3.5 6.5 42.25
Average = 3.5 713.5
Stock B
Year AR ER = Avg. R (AR-ER) (AR-ER)^2
2010 21 19.5 1.5 2.25
2011 33 19.5 13.5 182.25
2012 45 19.5 25.5 650.25
2013 -8 19.5 -27.5 756.25
2014 -7 19.5 -26.5 702.25
2015 33 19.5 13.5 182.25
Average = 19.5 2475.5
Average return of A= 3.5
Average return of B = 19.5
Volatility (Std Dev. Of A) = SQRT((∑(AR-ER)^2) / (n-1)) 11.94571
sQRT(713.5 / (6-1))
Volatility (Std Dev. Of B) = sQRT(2475.5 / (6-1)) 22.25084
(b)
Covariance between stock
Year Stock A (AR - ER) Stock B (AR-ER) Stock A (AR - ER) * Stock B (AR-ER)
2010 -16.5 1.5 -24.75
2011 16.5 13.5 222.75
2012 5.5 25.5 140.25
2013 -9.5 -27.5 261.25
2014 -2.5 -26.5 66.25
2015 6.5 13.5 87.75
753.5
Covariance = 753.5 / (6-1) 150.7
(c )
Correlation = Covariance between stock / (Std. dev. Of A * std. dev. Of B)
150.7 / (11.94571 * 22.25084)
0.566963
So, Correlation between two stock is 0.57.